### v1

# TAMO:Fine-Grained Root Cause Analysis via Tool-Assisted LLM Agent with Multi-Modality Observation Data

Abstract
With the development of distributed systems, microservices and cloud native technologies have become central to modern enterprise software development. Despite bringing significant advantages, these technologies also increase system complexity and operational challenges. Traditional root cause analysis (RCA) struggles to achieve automated fault response, heavily relying on manual intervention. In recent years, large language models (LLMs) have made breakthroughs in contextual inference and domain knowledge integration, providing new solutions for Artificial Intelligence for Operations (AIOps). However, Existing LLM-based approaches face three key challenges: text input constraints, dynamic service dependency hallucinations, and context window limitations. To address these issues, we propose a tool-assisted LLM agent with multi-modality observation data, namely TAMO, for fine-grained RCA. It unifies multi-modal observational data into time-aligned representations to extract consistent features and employs specialized root cause localization and fault classification tools for perceiving the contextual environment. This approach overcomes the limitations of LLM in handling real-time changing service dependencies and raw observational data and guides LLM to generate repair strategies aligned with system contexts by structuring key information into a prompt. Experimental results show that TAMO performs well in root cause analysis when dealing with public datasets characterized by heterogeneity and common fault types, demonstrating its effectiveness.
Index Terms:
Root cause analysis, Tool-Assisted LLM agent, Cloud-native systems, Multimodal data, Diffusion.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
In recent years, Microservice architecture and cloud-native technologies have become foundational elements in modern enterprise software development, driving the rapid development of distributed systems. [28],[32],[39]. Microservices decompose traditional monolithic applications into independent distributed services, while cloud native technologies leverage containerization and container orchestration[21] to enhance the capabilities for flexible resource allocation, dynamic scaling, and continuous delivery[30]. In enterprise production environments, microservice systems are composed of numerous independent services and containerized entities[45], which significantly increased system complexity, These systems have intricate inter-service dependencies and communication patterns, where the failure of any component (e.g., network resource corruption) can trigger cascading effects that jeopardize overall system stability, potentially leading to significant economic losses[2]. How to achieve accurate root cause identification and resolution in complex distributed systems has become a key factor in reducing production risks and minimizing losses for modern enterprises.
Figure 1: A failure case in a cloud-native system demonstrates three major challenges when using LLM directly for root cause analysis.
Root cause analysis aims to quickly correlate related multiple entity features, identify the root cause and determine the corresponding fault type, and provide actionable solutions, thus realizing an automatic fault response mechanism and minimizing human intervention[42]. How to perform an accurate and efficient root cause analysis in the distributed environment has become an important research direction. Traditional RCA methods typically rely on rule-based anomaly detection or statistical correlation analysis (e.g. causal graphs[24]) to localize faults. These methods struggle to jointly analyze heterogeneous data modalities (e.g., unstructured logs and structured metrics), often leading to fragmented information and frequent error detection. DL-based RCA methods attempt to incorporate multimodal feature fusion and dependency-sensitive Graph Neural Networks (GNNs)[23][44] to achieve accurate fault localization. Although these methods demonstrate good performance in specific tasks such as fault localization in complex systems, they still face significant challenges in real-time analysis and fault interpretability, heavily relying on the expertise of engineers to complete the root cause analysis and remediation process. Recently, large language models have demonstrated revolutionary capabilities in automated root cause analysis, opening new avenues for AIOps [9], [12]. LLM can combine system faults with the current system context to generate actionable fault diagnosis reports, as evidenced by prototypes such as RCACopilot[6]. Recent works such as RCAgent[35] and mABC[43] developed LLM-driven automated RCA frameworks that enable accurate fault localization and provide reasonable repair suggestions, thus further advancing the field of automated operations and maintenance.
However, although existing LLM-based methods have significantly alleviated the reliance on expert knowledge in traditional DL-based methods, they still face the following three key challenges in real distributed system environments:
  • Context Limitations: Limited by fixed-length context windows, the LLM is unable to effectively process massive amounts of contextual monitoring data in real-time. This constraint often leads to data truncation or compression during analysis, making it difficult to utilize historical information to detect anomalous trends in system components, thereby affecting the rationale and accuracy of root cause analysis. So how to extract useful information from the complex system context and input it into the context-limited LLM is a key challenge.
  • Text Input Constraints: Existing large language models such as ChatGPT[1], DeepSeek[27], Qwen[3], etc., either only support text data as input or enable modality input by mapping different modal features to a unified textual semantic space. However, due to the lack of obvious correlation between time series and text, this transformation process often results in a significant loss of key feature information, severely hindering the analytical and reasoning capabilities of LLM. So how to efficiently unify multiple modalities of observation data to assist the LLM in providing accurate root cause analysis results is a key challenge.
  • Dynamic Dependence Hallucination: The LLM is based on static, pre-trained knowledge, which makes it challenging to model service dependencies in systems. It can only simulate simple graph structures using text prompts. As noted in [34], LLMs struggle with reasoning over complex graphs, which limits their ability to effectively capture service dependencies. Additionally, the dynamic nature of service call chains in the system means that evolving dependencies can cause the LLM to generate hallucinations that contradict actual dependencies, leading to inaccurate root cause analysis. So how to assist the LLM in modeling fault propagation paths using call chain relationships to achieve fine-grained root cause localization is a key challenge.
To address these limitations, we propose a tool-assisted LLM agent framework based on domain-specific perceptive models. This framework decouples the LLM from the raw observational data by utilizing specialized observational data analysis tools($\mathscr{T}_{1}$-$\mathscr{T}_{3}$), addressing the constraints of text input and the limitations of context in the LLM. These specialized model tools are used to perceive the contextual environment, and their results form structured text inputs that are fed back to the RCA expert agent for further analysis, generating accurate root cause analysis results and reasonable maintenance recommendations. Specifically, to address the problem of multimodal data alignment, we propose a dual-branch diffusion model as the data feature observation tool ($\mathscr{T}_{1}$), which unifies multimodal observability data (logs, traces, metrics) into temporally consistent representations. To address the issue of dynamic dependencies, the root cause localization tool ($\mathscr{T}_{2}$) performs a joint analysis of entity features and call topologies, modeling the fault propagation chain, and identifying the root cause through frequency-domain attention-driven temporal causal analysis. Subsequently, the fault classification tool ($\mathscr{T}_{3}$) further analyzes the fault and maps the symptoms to predefined fault types. The main contributions are summarized as follows:
  • We propose a tool-assisted LLM agent framework, which integrates domain-specific model tools with LLMs for root cause analysis of microservices. This framework enables the LLM to analyze the system state in real time by calling perception tools, decoupling the agent from the raw observational data, allowing it to generate accurate root cause analysis results and reasonable maintenance recommendations.
  • We present a dual-branch collaborative diffusion tool that captures potentially consistent patterns across modalities and provides a unified time series representation for downstream analysis tasks through condition-oriented collaborative reconstruction.
  • We design a root cause localization tool based on frequency-domain attention mechanisms, which identifies causal dependency patterns in frequency-domain features while filtering out low-frequency noise, achieving precise root cause localization.
  • Experiments on two benchmark datasets demonstrate that TAMO outperforms all compared methods including state-of-the-art (SOTA) approaches and derived baselines. Ablation and case studies further validate the effectiveness of the framework.
II Related Work
In modern cloud native systems, with the widespread adoption of microservice architectures and dynamic infrastructures, it has become particularly important to ensure high system availability and rapid fault recovery, making root cause analysis (RCA) a critical part of operations and maintenance. As a result, various RCA methods have been developed to guarantee system reliability, including traditional RCA methods, multi-modal RCA methods, and LLM-based RCA methods.
Traditional RCA methods: Traditional RCA methods are mainly based on statistical correlation analysis techniques, such as anomaly detection based on predefined rules[37] and root cause localization strategies based on causal discovery[7]. While these approaches perform well in static environments, they have difficulties in adapting to the dynamic service dependencies that characterize cloud native environments. With the rapid advancement of deep learning, neural network-based methods for real-time system data analysis continue to emerge. DeepLog[11] models log features via long short-term memory networks; GDN[10] and GTA[8] construct graph neural networks using metrics data to capture fault propagation paths; Sage[13] leverages causal Bayesian networks with trace data for root cause localization. However, these approaches predominantly focus on unimodal data, failing to exploit complementary inter-modal relationships.
Multi-modal RCA methods: The multi-modal-based approach significantly reduces the missed detection rate by comprehensively considering system observation data from different modalities. Due to the differences between modalities, it is crucial to extract the correlated features across different data. PDiagnose[18] analyzes different modality data as independent entities and uses a voting mechanism to identify the root cause entity. Although this method successfully utilizes data from multiple modalities, it overlooks the consistency relationships between features from different modalities. Eadro[23], performs gated fusion of features extracted from different modalities and uses a graph neural network to fuse embedded features at the service level for root cause localization. HolisticRCA[14], compared to Eadro, fully considers the heterogeneous nature of microservice systems and standardizes the embedding of different modality data using an assembling building blocks strategy, combining masked embeddings for overall root cause analysis. In contrast to directly using raw data for root cause localization, Nezha[40] and DiagFusion[42] convert heterogeneous multimodal data into homogeneous event representations, performing joint analysis through event graphs to achieve root cause localization. CoE[38] further supports the operational experience of site reliability engineers as input on the event graph, enhancing the understanding and interpretability of the event. While these methods significantly improve the accuracy of root cause localization through multimodal fusion and comprehensive system analysis, their root cause analysis process still relies on human intervention by engineers.
LLM-based RCA methods: Due to the advanced reasoning and contextual understanding capabilities of large language models, LLM-based RCA methods have become significant advancements in automated root cause analysis. Currently, LLM-based RCA methods are still in the early stages of development. RCAgent[35] handles anomalies through a tool-augmented agent framework, while mABC[43] employs a blockchain-based multi-agent system for root cause voting and solution formulation. However, directly applying LLMs to RCA presents several critical challenges: the limited context window of LLMs restricts the processing of large-scale observation data, and due to text input constraints, they are unable to effectively handle time-series data (e.g., metrics), resulting in the loss of important feature information. Additionally, another key issue is how to use large models to capture and model dynamic dependencies in cloud native systems in real-time, without introducing misleading or false conclusions. These challenges highlight the need for further innovations in integrating LLM with multimodal and heterogeneous data sources for more accurate root cause analysis.
Figure 2: The proposed TAMO framework consists of a Data Feature Observation Tool ($\mathscr{T}_{1}$), a Root Cause Localization Tool ($\mathscr{T}_{2}$), a Fault Classification Tool ($\mathscr{T}_{3}$), and an RCA Expert Agent ($\mathscr{A}$). In the framework, the agent $\mathscr{A}$ calls tools $\mathscr{T}_{1}$-$\mathscr{T}_{3}$ as perception tools to analyze the system contextual observational data in real time. The results of these perceptions are structured into text inputs that are fed back into the agent. By analyzing the perception results, the expert agent will provide the corresponding root cause analysis and repair suggestions.
III METHODOLOGY
To fully exploit the multimodal observability data of cloud native systems while integrating prior domain knowledge, we propose TAMO, Tool-Assisted LLM Agents that Integrates Multi-Modality Observation Data to address root cause entity localization and fault classification in large-scale system operations and maintenance. This workflow utilizes observation tools composed of a dual-branch collaborative diffusion model to effectively extract and integrate multimodal data from different entities. Additionally, task-specialized tools explicitly model the complex dependencies among entities using a causal graph, with feature propagation through the causal graph enabling accurate root cause analysis. Furthermore, to fully incorporate domain knowledge, we employ a Large Language Model (LLM) with operational expertise as an expert agent. The expert agent $\mathscr{A}$ integrates the outputs from the preceding tools $\mathscr{T}_{2}$-$\mathscr{T}_{3}$ with context knowledge of the system to offer human engineers a comprehensive fault analysis and recommended solutions. As shown in Fig.2, the proposed workflow consists of three tools and one agent, $\mathscr{T}_{1}$$\mathscr{T}_{3}$ and $\mathscr{A}$, each responsible for a specific task in the root cause analysis process:
  1. 1.
    Data Feature Observation Tool ($\mathscr{T}_{1}$): This tool employs a dual-branch collaborative diffusion model to integrate multimodal features. It reconstructs and fuses textual data features, such as logs, into time-series data based on conditional guidance;
  2. 2.
    Root Cause Localization Tool ($\mathscr{T}_{2}$): Utilizing the multimodal-enhanced time-series data as input, this tool models causal relationships between entities using a frequency-domain self-attention module, enabling accurate fault root cause localization;
  3. 3.
    Fault Classification Tool ($\mathscr{T}_{3}$): Based on the causal relationships between entities and the invocation chain structure, this tool classifies the types of faults;
  4. 4.
    RCA Expert Agent ($\mathscr{A}$): Leveraging the expertise of a Large Language Model (LLM), this agent synthesizes the outputs from previous tools along with system environment context to provide site reliability engineers (SREs) with a comprehensive fault analysis and recommended solutions.
III-A Data Feature Observation Tool ($\mathscr{T}_{1}$)
Diffusion models are a class of generative models based on Markov processes that have demonstrated exceptional performance in the generation of data in various domains, including images[19], voice[31], and time series[41]. The core idea is to gradually add noise to the data through a forward denoising (forward diffusion) process, which converts the original data distribution to an approximate Gaussian distribution. In the reverse denoising process, the model learns how to gradually remove the noise to generate the desired data. In this tool, we use log patterns and temporal features as control conditions for two diffusion branches. During the denoising process, the two branches collaborate to guide the generation of time-series data with multimodal features from the noise-perturbed original data. Specifically, during the forward diffusion process, the original data $x_{0}\sim q(x_{0})\in\mathbb{R}^{E\times m\times l}$ consists of $E$ entities, $m$ channels, and a time series of length $l$. This process can be expressed as:
$q\left(x_{1},\cdots,x_{T}\mid x_{0}\right):=\prod_{t=1}^{T}q\left(x_{t}\mid x_{t-1}\right),$ (1)
If $\alpha_{t}\in(0,1)$ is a predefined noise scaling factors that determines how much information from the previous step is retained, then each step of the Eq.1 is:
$q\left(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}\right)=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{t-1},\left(1-\alpha_{t}\right)\mathbf{I}\right),$ (2)
After $T$ steps of noise addition, $x_{T}$ converges to random noise $x_{T}\sim\mathcal{N}(0,I)$, which approximates an Gaussian distribution.
Subsequently, we perform stepwise denoising on the noisy data using a dual-branch conditional mechanism. The parameterized conditional transition distribution can be expressed as:
$p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathcal{C}\right)$ $=\mathcal{N}\left(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t,\mathcal{C}\right),\mathbf{\Sigma}_{\theta}\left(\mathbf{x}_{t},t,\mathcal{C}\right)\right),$ (3)
$p_{\theta}\left(\mathbf{x}_{0:T},\mathcal{C}\right)$ $=p\left(\mathbf{x}_{T}\right)\prod_{t=1}^{T}p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathcal{C}\right),$
where $\mathcal{C}=\{c_{log},c_{time}\}$ represents the set of control conditions, with $c_{log}$ and $c_{time}$ corresponding to the log branch and the time-series branch, respectively. These conditions guide the denoising process, ensuring that the generated data have multi-modal features by utilizing both log-based patterns and temporal features.
For the generation of the log-conditioned variable $c_{log}$, since unstructured log representations are difficult to extract high-quality features, inspired by previous work [14] [23], we extract event patterns $\mathcal{E}$ and keywords $\mathcal{W}$ from the raw logs using Drain[16], and by calculating the TF-IDF scores, we select the highest scoring pattern-keyword pairs as the conditional variable $c_{log}$.
$c_{log}=\left\{\arg\max_{e\in\mathcal{E}}\operatorname{TF-IDF}(e)\right\}\cup\left\{\arg\max_{w\in\mathcal{W}}\operatorname{TF-IDF}(w)\right\},$ (4)
For the generation of the time-conditioned variable $c_{time}$, we extract key performance indicators (KPIs) such as latency and duration from trace files. These KPIs $x_{kpi}$ are then embedded into a latent space using a Multi-Layer Perceptron (MLP) network, enabling the model to effectively incorporate temporal characteristics into the denoising process.
$\mathbf{c}_{time}=\operatorname{Embedding}\left(x_{kpi}\right)\in\mathbb{R}^{h_{{time}}},$ (5)
During training, we use a dual-branch network to learn the reverse diffusion process, gradually denoising the noisy data. In this process, the noise term estimated at each step is collaboratively generated by the two branches.
$\mathbf{X}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{X}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}\left(\mathbf{X}_{t},t,\mathcal{C}\right)\right)+\sigma_{t}\mathbf{z}$ (6)
$\mathbf{z}\sim\mathcal{N}(0,\mathbf{I}),$
in which, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ represents the cumulative product of the noise scaling factors $\alpha_{s}$ and $\epsilon_{\theta}$ represents the noise term predicted by the model, which is jointly generated based on the computations from the log branch $\epsilon_{log}$ and time-series branch $\epsilon_{time}$. A hyperparameter $\mu$ is used as a mixing coefficient to balance the contributions of the two branches.
$\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathcal{C}\right)=\mu\epsilon_{log}\left(\mathbf{x}_{t},t,\mathbf{c}_{log}\right)+\left(1-\mu\right)\epsilon_{time}\left(\mathbf{x}_{t},t,\mathbf{c}_{time}\right),$ (7)
The original denoising diffusion model employs U-Net as the denoising backbone [17]. However, considering the characteristics of time-series data, we adopt a patch-based Transformer network [29] as the backbone for each branch to better capture the representations of time-series subsequences. The training objective of this model is to minimize the KL-divergence between the distributions $p$ and $q$. According to [17], the objective derivation is formulated as follows:
$\mathcal{L}_{diff}=\mathbb{E}_{\mathbf{x}_{0},\boldsymbol{\epsilon}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(x_{t},t,\mathcal{C})\right\|_{2}^{2}\right],$ (8)
III-B Root Cause Localization Tool ($\mathscr{T}_{2}$)
Through the dual-branch diffusion model of Tool $\mathscr{T}_{1}$, we obtain time-series data that integrates multimodal features. However, since this process operates at the level of system resource entities, the generated time-series $x^{\prime}_{ent}\in\mathbb{R}^{L\times d_{ent}}$ maintains the same dimensionality as the original resource entities, where $ent$ represents the corresponding resource entity, $L$ denotes the time window length, and $d_{ent}$ represents the feature dimension associated with each entity. Due to the heterogeneity of observability data across different resource entities in cloud-native systems, inconsistencies exist among them, which hinders unified learning in the model. To address this issue, we first perform data embedding on the features of different entities, mapping them into a unified dimensional space for consistency:
$H^{ent}=\operatorname{DataEmbedding}\left(x^{\prime}_{ent}\right)\in\mathbb{R}^{L\times d_{model}},$ (9)
where $d_{model}$ represents the unified embedding dimension. After mapping all entity features into the latent space, we concatenate the output results to obtain the embedded vector $H^{emb}\in\mathbb{R}^{E\times L\times d_{model}}$.
After obtaining the unified embedding vector $H^{emb}$, we consider that root cause localization tasks typically require the detection and identification of short-term anomalies or disturbance signals within the system entity. These transient features may be obscured by smoothed trends in the time domain. To address this issue, we apply the Fast Fourier Transform (FFT) to convert the data from the time domain to the frequency domain. Formally, the transformation can be expressed as:
$\mathcal{F}\left(H^{emb}\right)=\operatorname{FFT}\left(H^{emb}\right)=\sum_{i=0}^{L-1}H^{emb}_{i}e^{-j2\pi ki/K},\quad$ (10)
where $e^{-j2\pi ki/N}$ is the complex exponential Fourier basis function, $H^{emb}_{i}$ represents the time-series value corresponding to the time step $i$, and $k=0,1,\ldots,N-1$.
Considering the real-time demands of root cause localization, the model needs to focus on short-term fluctuations in the observed data. To achieve this, we apply a high-pass filter to the frequency-domain data, retaining the first $k$ high-frequency components which helps to amplify the anomalous signals within the data. Formally, this can be expressed as:
$\operatorname{Indices}=\operatorname{argtopk}\left(\left|\mathcal{F}\left(H^{emb}\right)\right|,k=freq\_topk\right),$ (11)
$Mask=\begin{cases}1,&\text{ if }k\in\text{ Indices }\\ 0,&\text{ otherwise }\end{cases},k=0,1,\ldots,K-1,$ (12)
$\mathcal{F}_{\text{filtered }}=\mathcal{F}\left(H^{emb}\right)\odot Mask\in\mathbb{R}^{E\times K\times d_{model}},$ (13)
Since the interactions between microservices in the cloud native system can be described by a dependency graph, in order to fully consider the relationships between entities in the system and capture potential fault propagation pathways, we use the self-attention mechanism to construct a causal dependency graph in the frequency domain. The process can be defined as:
$A=Softmax\left(Attention(\mathcal{F}_{\text{filtered }}W^{Q},\mathcal{F}_{\text{filtered }}W^{K})\right)$ (14)
where $W^{Q}$,$W^{K}$ are the query and key matrices respectively. Subsequently, to reduce the influence of irrelevant entities and remove redundant relationships, we introduce topological constraints into the graph generation process. These constraints can be defined as:
$\mathbf{A}_{\text{maske}}=\mathbf{A}\odot\left(\mathbf{I}-\mathbf{I}_{E}\right)\odot\operatorname{TopkMask}(\cdot,k=topk)$ (15)
in which $\mathbf{I}_{E}$ is the entity diagonal matrix for avoiding self-loops and $\operatorname{TopkMask}$ preserves the top $k$ most significant dependencies for each entity. Then, we apply graph neural network to perform root cause inference and localization. Specifically, we select the Graph Attention Network (GAT) [4] as the backbone for aggregating neighborhood feature information among system entities, in order to simulate the fault propagation process as follows:
$\mathcal{H}^{s+1}_{i}=Relu\left(\alpha_{ii}W\mathcal{H}^{s}_{j}+\sum_{j\in N(i)}\alpha_{ij}W\mathcal{H}^{s}_{j}\right),$ (16)
where $N(i)=\{j|A_{mask}[i,j]=1\}$, $\mathcal{H}^{s}_{i}$ is the representation of entity $i$ in the $s$-th layer and $\mathcal{H}^{0}_{i}=\mathcal{F}_{\text{filtered }}$. $\alpha_{i,j}$ is the attention coefficient between system entity $i$ and its related entity $j$, which can be calculated by the following formula:
$\xi(i,j)=LeakyRelu\left(a^{T}\left(W^{s}\cdot h^{s}_{i}\oplus W^{s}\cdot h^{s}_{j}\right)\right),$ (17)
$\alpha_{ij}=\frac{\operatorname{exp}(\xi(i,j))}{\sum_{p\in N_{i}\cup\{i\}}\operatorname{exp}(\xi(i,p))},$ (18)
where $\oplus$ denotes the concatenation operation, $a^{T}$ represents the vector of attention parameters and $W^{s}$ is the weight matrix for the $s$-th layer. Finally, we use a max-pooling layer to aggregate the importance of instances, thereby obtaining the root cause probability for the system entities:
$\hat{y}=\text{ MaxPool}\left(\mathbf{H}^{S}\right)\in\mathbb{R}^{E\times 1},$ (19)
Considering that multiple types of entities may experience faults simultaneously within the same time period, meaning that the root cause entity may not be unique, we use binary cross-entropy (BCE) as the loss function to optimize the model. Specifically, for each label $y_{i}$ and the model predicted value $\hat{y_{i}}$ (where $i\in\{0,1,...,E-1\}$, and $E$ represent the total number of system entities) , the loss function is defined as:
$\mathcal{L}_{BCE}=-\frac{1}{E}\sum_{i=1}^{E}\left[y_{i}\log(\hat{y_{i}})+(1-y_{i})\log(1-\hat{y_{i}})\right],$ (20)
III-C Fault Classification Tool ($\mathscr{T}_{3}$)
After the root cause entity localization is performed by Tool $\mathscr{T}_{2}$, the fault diagnosis process typically requires further identification of the specific fault type of the root cause entity (e.g., hardware overload, resource leakage, or configuration errors) to guide precise remediation strategies. Although root cause localization narrows down the scope of fault analysis (i.e., determining “where” the anomaly occurs), in practical operational scenarios, different fault types often correspond to differentiated handling strategies (e.g., memory leaks require service restarts, while CPU contention necessitates resource scaling). Therefore, this tool uses the time series data $x^{\prime}_{ent}$ generated by tool $\mathscr{T}_{1}$ (for simplicity, denoted as $z$) to train a fault classification model.
To model the latent temporal dependencies in the input time-series data $z$, we use a Transformer model [33] as the encoder network. The output of its single-head self-attention mechanism is given by:
$\text{Attention}=\text{softmax}\left(\frac{zW^{Q}\cdot z{W^{K}}^{T}}{\sqrt{d_{k}}}\right)zW^{V},$ (21)
where $W^{Q}$,$W^{K}$,$W^{V}$ are linear projection matrices. Subsequently, we concatenate the results of the multi-head attention to obtain the latent vector representation:
$V=\text{MultiHead}\left(zW^{Q},zW^{K},zW^{V}\right),$ (22)
After capturing the internal temporal dependencies of the time-series data, we employ the GAT model to capture the spatial dependencies between the time steps. According to formulas Eq.16-Eq.18, we perform feature propagation on the latent vector representation. The representation of each entity is obtained by aggregating the weighted features of its neighboring entities:
$Z^{\prime}=\text{ELU}\left(\text{GATConv}(X,A)\right),$ (23)
where ELU is the activation function and $A\in\mathbb{R}^{E\times E}$ is the dependency between entities. Subsequently, we use an MLP network to output the fault type for the corresponding entity. Considering the issues of multi-class overlap and class imbalance in the fault classification task, we adopt a weighted binary cross-entropy (wBCE) loss function with the introduction of class weight coefficients $\lambda$ as the model’s training objective:
$\mathcal{L}_{wBCE}=-\frac{1}{E}\sum_{i=1}^{E}\left[\lambda_{w}\cdot y_{i}\log(\hat{y_{i}})+(1-y_{i})\log(1-\hat{y_{i}})\right],$ (24)
in which $\lambda_{w}$ represents the ratio of negative to positive samples within each fault type class, which is used to enhance the loss contribution of the samples. Additionally, to prevent overfitting, we include an L2 weight decay term in the loss function. The final loss function can thus be formulated as:
$\mathcal{L}=\mathcal{L}_{wBCE}+\beta\left||\mathbf{W}\right||^{2}_{2}.$ (25)
where $\beta$ is a predefined hyperparameter and $\mathbf{W}$ is the model weights.
III-D RCA Expert Agent ($\mathscr{A}$)
After the three tools $\mathscr{T}_{1}$-$\mathscr{T}_{3}$ have completed tasks such as data feature alignment, root cause localization, and fault classification, the workflow has achieved a relatively deep understanding of the fault. However, relying solely on these outputs makes it difficult to provide comprehensive and accurate failure analysis and remediation plans. Therefore, to further improve fault response and system recovery efficiency, we introduce the RCA expert agent $\mathscr{A}$. Using large language models (LLM), this agent synthesizes the predictions from the previous tools and incorporates the current context of cloud-native microservice architecture to offer a more comprehensive and accurate fault diagnosis and resolution strategy for site reliability engineers.
Specifically, we use the GPT-4 [1] as the expert model, which gathers the outputs from the previous tools as background knowledge, including:
  • Root Cause Localization Results (from $\mathscr{T}_{2}$): Identifying the root cause location of the fault and the associated system components.
  • Fault Classification Results (from $\mathscr{T}_{3}$): Providing the specific type of fault, such as network failure, hardware failure, configuration errors, etc.
  • System Context Information: Including the system microservice architecture, dependencies between services, call chains, resource usage, and other related details, which provide essential context for fault diagnosis.
RCA Expert Prompt You are an RCA expert responsible for analyzing and diagnosing failures of complex systems based on cloud-native, microservices architectures. You have access to detailed failure reports, system performance data, and historical failure patterns. You provide detailed analysis based on the following structured inputs, which include root cause information, failure classification, system architecture context, historical failure data, and other system state details. You are expected to apply advanced reasoning methods, drawing on knowledge of system architecture, historical failure modes, and industry best practices, to provide comprehensive, practical solutions to improve system reliability and minimize downtime. [Failure root cause localization] - Root Entities: <localization results from $\mathscr{T}_{2}$> - Related Entities: <the related calling entities> [Fault Classification] - Fault Type: <classification results from $\mathscr{T}_{3}$> - Symptoms: <detailed Symptom Description> [System architecture background] - Architecture: <microservice functionality> - Context: <related logs and trace information> [Task Objective] Based on the information above, please analyze the root cause, impact scope, and proposed solutions for the fault, and provide the following: 1. Analysis of the root cause of the fault. 2. Analysis of the type and level of failure. 3. Repair solutions and preventive measures.
Figure 3: Prompt templates for this RCA expert agent.
In practice, we prepare the prompt in the format shown in Fig.3 as input to the model, and the expected output is a detailed fault analysis report along with corresponding solution recommendations. Such outputs assist the SRE team in quickly understanding the root cause, impact scope, and taking targeted remediation actions, thus enhancing system reliability and response efficiency.
TABLE I: The experimental results of different approaches on the tasks of root cause localization and fault classification.
Data Approach RE Localization Fault Type Classification
AC@1 AC@3 AC@5 MiPr MaPr MiRe MaRe MiF1 MaF1
$A_{s}$ HolisticRCA 65.62% 76.33% 81.69% 0.6355 0.7130 0.7728 0.8037 0.6974 0.7507
TimesNet 14.81% 32.10% 61.72% - - - - - -
Eadro 13.58% 37.03% 45.68 % 0.1887 0.0770 0.1235 0.1046 0.1493 0.0758
SegRNN 14.81% 40.75% 60.49% - - - - - -
DejaVu 62.50% 65.62% 78.13% - - - - - -
LightGBM - - - 0.1967 0.1536 0.2000 0.2282 0.1983 0.1707
Transformer 16.22% 37.84% 56.76% 0.0114 0.0123 0.1892 0.1932 0.0214 0.0230
TAMO 71.87% 82.14% 88.83% 0.7164 0.6671 0.6486 0.6149 0.6809 0.6299
$A_{p}$ HolisticRCA 65.62% 80.80% 86.60% 0.5913 0.6682 0.7534 0.7598 0.6626 0.7008
TimesNet 10.71% 22.14% 32.14% - - - - - -
Eadro 9.38% 13.13% 15.63% 0.3901 0.3344 0.4665 0.4792 0.4249 0.3761
SegRNN 18.75% 56.25% 84.38% - - - - - -
DejaVu 18.75% 21.88% 28.13% - - - - - -
LightGBM - - - 0.6849 0.6450 0.6792 0.6808 0.6820 0.6583
Transformer 10.71% 19.29% 24.29% 0.0106 0.0106 0.4254 0.4439 0.0207 0.0208
TAMO 64.28% 80.36% 87.50% 0.7182 0.6597 0.6840 0.6642 0.7007 0.6445
$A_{n}$ HolisticRCA 73.66% 75.89% 79.01% 0.6219 0.6437 0.7612 0.7322 0.6846 0.6717
TimesNet 21.15% 48.07% 78.85% - - - - - -
Eadro 17.18% 28.13% 42.19% 0.5426 0.6003 0.7969 0.8145 0.6456 0.6575
SegRNN 7.50% 20.63% 23.75% - - - - - -
DejaVu 81.25% 90.63% 100.00% - - - - - -
LightGBM - - - 0.5208 0.5491 0.8333 0.8178 0.6410 0.6406
Transformer 13.46% 48.08% 78.85% 0.1415 0.1423 0.5577 0.5207 0.2257 0.2212
TAMO 84.37% 91.96% 97.77% 0.8718 0.8869 0.8947 0.8681 0.8831 0.8706
$B$ HolisticRCA 61.11% 75.00% 77.78% 0.5000 0.5936 0.8056 0.8056 0.6170 0.6636
Timesnet 15.62% 25.00% 40.63% 0.4893 0.5789 0.7187 0.7398 0.5823 0.6394
Eadro 40.62% 56.25% 71.87% 0.4167 0.5211 0.6250 0.6516 0.5000 0.5694
SegRNN 18.75% 43.75% 62.50% 0.1496 0.2359 0.5937 0.6287 0.2389 0.3296
DejaVu 15.62% 34.38% 50.00% - - - - - -
LightGBM - - - 0.7832 0.6434 0.4186 0.3860 0.4687 0.4122
Transformer 46.88% 68.75% 84.38% 0.4464 0.5664 0.7813 0.7955 0.5682 0.6406
TAMO 72.22% 80.55% 86.11% 0.8500 0.8750 0.5313 0.5682 0.6538 0.6421
IV Evaluation
IV-A Research Questions
This section answers the following research questions:
  • RQ1: How effective is TAMO in root cause localization?
  • RQ2: How effective is TAMO in Fault Classification?
  • RQ3: How well does TAMO work when ablation studies are performed on its various components?
  • RQ4: How well does TAMO perform in real-world case studies and fault scenarios?
IV-B Experiment Setup
IV-B1 Datasets
We conducted extensive experiments on two public datasets:
  • Dataset $A$. $A$ is a large-scale public dataset from the AIOps Challenge [^1] https://competition.AIOps-challenge.com/home/competition/1496398526429724760, which was collected through fault injection into the real-world deployed microservice system HipsterShop [^2] https://github.com/GoogleCloudPlatform/microservices-demo. The dataset primarily consists of three types of data: metrics, logs, and traces. The system uses a dynamic deployment architecture, consisting of 10 services, each with 4 pods, totaling 40 pods that are dynamically deployed across 6 nodes. The dataset includes 15 types of faults in total. Among these, 9 fault types are related to services and pods in the context of Kubernetes (K8s) containers [5], while node faults include 6 types: sudden memory pressure, disk space exhaustion, disk read/write issues, CPU pressure, and slow CPU growth.
  • Dataset $B$. $B$ is a public dataset [^3] https://doi.org/10.5281/zenodo.7615393 collected from the broadcast-style SocialNetwork system in [23], which only contains service-level entities. It includes 21 microservices and also consists of three types of data: metrics, logs, and traces. The dataset was generated by injecting faults into the system using Chaosblade, producing data that includes three types of faults: CPU resource exhaustion, network latency, and packet loss.
IV-B2 Baselines
We compared TAMO with the most popular integrated multimodal methods and derivative methods from the time series domain. Specifically, we selected two state-of-the-art integrated multimodal root cause analysis methods, Eadro[23] and HolisticRCA[14], as well as four unimodal time series analysis methods (i.e., TimesNet[36], SegRNN[26], DejaVu[25], Transformer[33]). These methods were implemented using their publicly available and reproducible open-source code, and we adjusted the data format to match the required input. Furthermore, since our research task is relatively novel and includes both fault localization and classification, we incorporated the same fault classifier used in TAMO for methods that do not natively implement fault classification (such as methods [36], [23], [33], etc.) for comparison. Furthermore, we introduced LightGBM[20], a method specifically designed for fault classification tasks, as a benchmark to evaluate TAMO performance.
TABLE II: The experimental results of the ablation study conducted on our proposed TAMO framework.
Models RE Localization Fault Type Classification
AC@1 AC@3 AC@5 MiPr MaPr MiRe MaRe MiF1 MaF1
$\operatorname{TAMO}_{w/o\;\mathscr{T}_{1}}$ 43.75% 56.25% 68.75% 0.5909 0.5556 0.4063 0.4453 0.4815 0.4911
$\operatorname{TAMO}_{w/o\;branch}$ 59.38% 68.75% 71.87% 0.7200 0.7350 0.5625 0.6010 0.6316 0.6167
$\operatorname{TAMO}_{w/o\;FFT}$ 53.13% 68.75% 78.13% 0.8000 0.7639 0.5000 0.5404 0.6154 0.5977
$\operatorname{TAMO}$ 72.22% 80.55% 86.11% 0.8500 0.8750 0.5313 0.5682 0.6538 0.6421
IV-B3 Evaluation metrics
TAMO is dedicated to locating root cause instances and identifying fault types. Based on previous research works[14],[23],[42],[15], we selected different evaluation metrics to evaluate the performance of model for these two tasks. Specifically, for the root cause localization task, we use top-k accuracy (AC@K) as the evaluation metric, which represents the probability that the true root cause entity appears within the top-k results produced by the method. Here, we set $k=\{1,3,5\}$. Given $S$ as the set of faults, $RC$ is the ground truth root cause, and $rc[k]$ are the top-k root causes generated by the model. Therefore, AC@K can be defined as follows:
$AC@k=\frac{1}{|S|}\sum_{i\in S}^{S}\begin{cases}1,&\text{if }RC_{i}\in rc_{i}[k]\\ 0,&\text{otherwise}\end{cases}$ (26)
For the fault classification task, we use $Precision(Pre)$, $Recall(Rec)$, and $F1-score(F1)$ to evaluate performance. Given true positives (TP), false positives (FP), and false negatives (FN), the formulas for these metrics are as follows: $Pre=\frac{TP}{TP+FP}$,$Rec=\frac{TP}{TP+FN}$, $F1=\frac{2\cdot Pre\cdot Rec}{Pre+Rec}$. Considering that this task is a multi-label classification problem, we adopt both micro and macro metrics for a comprehensive evaluation. Micro metrics aggregate the performance across all labels, treating all instances equally, which is useful when the class distribution is imbalanced. In contrast, macro metrics compute the metric for each label independently and then average the results, giving equal weight to each label regardless of its frequency. Thus, the final performance is measured using micro-precision ($MiPr$), macro-precision ($MaPr$), micro-recall ($MiRe$), macro-recall ($MaRe$), micro-F1 score ($MiF1$), and macro-F1 score ($MaF1$).
IV-B4 Implementation
We implemented TAMO using Python torch 2.4.0 and CUDA 12.1. The experiments were conducted on a Linux server equipped with an NVIDIA GeForce RTX 3090 GPU, using Python 3.8. The model was trained with the Adam optimizer [22], and all experiments used a fixed random seed. For hyperparameters, we set the batch size to 32, $\mu$ to 0.5, $\beta$ to 0.001, and the initial learning rate to 0.001. Additionally, due to the heterogeneity of system entities in Dataset $A$, which includes three levels: Pod, Service, and Node, the collected data features vary across these levels, making it challenging to train certain baselines uniformly. For such baselines, we partition Dataset $A$ into subsets composed of different entity types (e.g., $A_{p}$, $A_{s}$, and $A_{n}$) and perform separate training and testing for each subset to evaluate their performance.
IV-C RQ1: Effectiveness in Root Cause Localization
Table I presents the results for root cause localization. The experimental data shows that our proposed TAMO model achieves the best or second-best performance across all datasets. Specifically, on the $A_{s}$ dataset, the score of TAMO exceeds the second-best score by at least 6%, and even in the second-best performance case, the gap with the best result is kept under 3%, clearly outperforming other competitors. Furthermore, we observe that the Eadro and Transformer models, which perform well on the B dataset, experience a significant drop in performance on heterogeneous datasets such as $A_{s}$, $A_{p}$, and $A_{n}$. This is because these two methods can only handle homogeneous entity features (e.g., in the dataset B, there is only one type of service entity). When confronted with heterogeneous entity features that need to be processed separately, their performance significantly deteriorates due to the lack of inter-entity correlation information. In contrast, both HolisticRCA and our approach model heterogeneous entities in a unified manner, maintaining strong performance even on the dataset A. Compared to HolisticRCA, we map multimodal information into time-series features and further optimize it through spatiotemporal modeling, achieving better performance in the root cause localization task.
IV-D RQ2: Effectiveness in Fault Classification
In the fault classification task, as shown in Table 2, TAMO demonstrates significant advantages in two key metrics: Micro Precision (MiPr) and Macro F1 (MaF1). Particularly on the $A_{n}$ dataset, TAMO achieves MaPr of 0.8869 and MaF1 of 0.8706, marking improvements of 24.32% and 19.85% respectively over the second-best model(HolisticRCA). TAMO stands out across all metrics such as Precision, Recall, and F1. Furthermore, it exhibits remarkable performance in the classification of complex fault types at the node level. In addition, TAMO excels in the precision metric, with MiPr reaching 0.7164 and 0.7182 on $A_{s}$ and $A_{p}$, respectively, surpassing the best baseline model by 8. 09% and 12. 69%, significantly reducing the risk of misclassification. Although MiPr on the smaller dataset $B$ is lower than the baseline model LightGBM, TAMO still outperforms LightGBM in terms of MaPr and F1 metrics. This suggests that TAMO performs more evenly across different classes, achieving better classification results for minority classes with fewer samples.
Figure 4: In case study experiments, two types of faults were induced in microservices, with the root cause entities identified as the Currency Service and CheckoutPod0.
IV-E RQ3: Ablation Study
To explore the effectiveness of each component in our TAMO framework, we designed three variant models and conducted ablation experiments. The variant models are as follows:
  • $\operatorname{TAMO}_{w/o\;\mathscr{T}_{1}}$: To investigate the effectiveness of the diffusion method used by tool $\mathscr{T}_{1}$ in fusing multimodal feature, in this variant we removed the tool $\mathscr{T}_{1}$ and directly used the raw time series data as input to evaluate the variant performance.
  • $\operatorname{TAMO}_{w/o\;branch}$: To investigate the effectiveness of the dual-branch diffusion model in tool $\mathscr{T}_{1}$, we removed the branch $\epsilon_{time}$, which is responsible for maintaining temporal features, in this variant. Only the log branch $\epsilon_{log}$ was used in the diffusion reconstruction process. The reconstructed data were then used for fault classification to evaluate variant performance.
  • $\operatorname{TAMO}_{w/o\;FFT}$: To investigate the effectiveness of the frequency domain-based root cause localization method proposed for tool $\mathscr{T}_{2}$, we removed the Fourier process in this variant. Instead, the original time-domain data were directly used for subsequent causal graph generation and graph convolution processes. The root cause localization results were then output to evaluate variant performance.
Figure 5: Comparison of LLM results between RCA analysis of Agent $\mathscr{A}$ enhanced with TAMO tools and root cause analysis performed directly on raw data
Table II presents the results of the ablation experiments, from which it is evident that the removal of any component in TAMO leads to performance degradation. This demonstrates that each of the component within TAMO has an important effect. In particular, the variant of TAMO without tool $\mathscr{T}_{1}$ shows a significant decrease in performance. This is because the fact that the variant $\operatorname{TAMO}_{w/o\;\mathscr{T}_{1}}$ lacks the ability of the diffusion model to fuse multimodal information, thereby being reduced to learning with only a single modal feature. This absence of key feature information makes accurate root cause analysis challenging. Similarly, removing the time branch $\epsilon_{time}$ led to noticeable decreases in performance for both tasks, indicating that time series features are crucial for accurate fault localization and classification. Eliminating the FFT process resulted in a substantial degradation in RE localization metrics, demonstrating that frequency domain analysis is essential for precise root cause localization.
IV-F RQ4: Case Study
To validate the root cause analysis capabilities of TAMO in cloud-native ecosystems, this study conducts a case investigation using real-world failure data from HipsterShop. As shown in Fig. 4, the microservices system is deployed using a Kubernetes (K8s) architecture, with service components developed in multiple programming languages. The system entry is managed through the Frontend service, and each service node is configured with three Pod replicas. Monitoring data collection points are set at both the service and Pod levels. Experimental data shows that TAMO demonstrates precise root cause identification for the current failure scenario: the root cause localization confidence for the Currency Service reaches 97.96% (service level), while the confidence for CheckoutPod0 is 82.39% (Pod level). Notably, while the overall Checkout service is functioning correctly, only Pod0 experiences issues. In contrast, the Currency service suffers from a global failure at the service level, causing anomalies in Pods 0-2. This comparative case indicates that TAMO effectively integrates multi-entity features and topological causal relationships to achieve precise root cause diagnosis across multiple layers. In the fault classification task, we perform a visual analysis of observation data for the anomalous services/Pods: CheckoutPod0 is accurately classified as a CPU overload fault due to a significant spike in core metrics such as ”CPU Usage Second.” The Currency service exhibits a correlation between increased service-level response time and decreased request success rate, while its Pod-level Receive Packet count continues to decline. TAMO successfully identifies the anomaly as a network packet loss fault through multi-dimensional time-series pattern analysis.
Subsequently, we input the model inference results along with system context information into the RCA expert LLM and compare its output with that of an LLM fed directly with raw data, as shown in Fig.5. This framework leverages tool information and domain expertise to accurately analyze current system faults and provide reasonable remediation suggestions. In contrast, the normal LLM using only raw data struggles to handle the vast amount of observational data due to severe context limitations, resulting in significant omissions and misjudgments in root cause identification and fault classification, as well as corresponding errors in the proposed remediation suggestions.
V Conclusion
This paper introduces an innovative Tool-Assisted LLM Agent framework named TAMO to address the complex challenges of root cause analysis in cloud-native microservice environments. By integrating domain-specific tools with large language models, TAMO effectively handles high-dimensional, multi-modal cloud-native observability data and accurately pinpoints fault origins amidst dynamically changing service dependencies. The approach not only overcomes the information loss issues when converting raw observability data into a format suitable for LLM processing but also addresses the challenge that static pre-trained knowledge struggles to adapt to real-time changes in service dependencies. Moreover, by structuring and distilling key results, it guides the LLM to generate contextually appropriate remediation strategies, thereby avoiding limitations imposed by fixed-length context windows. Our research demonstrates that using a dual-branch collaborative diffusion model enables capturing consistent patterns across different data modalities, providing a unified temporal representation for downstream analysis tasks. Additionally, the root cause localization method based on frequency-domain attention mechanisms enhances the accuracy of identifying causal dependency patterns while effectively filtering out low-frequency noise, achieving precise fault localization. The experimental results show that, compared to existing advanced methods and benchmarks, TAMO performs notably better on two standard datasets, highlighting its potential to improve fault diagnosis efficiency and accuracy. Further ablation studies and case analyses also validate the effectiveness of each component of the framework and the superior performance of their collaborative work. However, as system scale and complexity continue to grow, how to more efficiently process larger volumes of data streams and further enhance real-time interpretability and cross-modal semantic alignment remain critical areas for future research.